Consider a project to supply Detroit with 31,000 tons of machine screws annually for automobile production. You will need an initial $6,200,000 investment in threading equipment to get the project started; the project will last for 5 years. The accounting department estimates that annual fixed costs will be $1,500,000 and that variable costs should be $285 per ton; accounting will depreciate the initial fixed asset investment straight-line to zero over the 5-year project life. It also estimates a salvage value of $875,000 after dismantling costs. The marketing department estimates that the automakers will let the contract at a selling price of $404 per ton. The engineering department estimates you will need an initial net working capital investment of $600,000. You require a return of 13 percent and face a tax rate of 24 percent on this project.

  

a-1.

What is the estimated OCF for this project? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

a-2. What is the estimated NPV for this project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
b. Suppose you believe that the accounting department’s initial cost and salvage value projections are accurate only to within ±15 percent; the marketing department’s price estimate is accurate only to within ±10 percent; and the engineering department’s net working capital estimate is accurate only to within ±5 percent. What are your worst-case and best-case NPVs for this project? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
Answer #1

a. Estimated OCF using tax shield approach

[{(404-285)*30000}-1500000]*0.76 + 0.24*(6200000/5)

1573200+2976000 = $1870800

b. NPV

-$6200000-$600000+$1870800(PVIFA13%,5 ) + [$600000+$875000(1-0.24)]/1.135

-$6200000-$600000+$1870800(3.5172) + [$600000+$875000(1-0.24)]/1.135

NPV= $466582.18

c. Worst case -

OCF = [{(404(1-10%)-285)*30000}-1500000]*0.76 + 0.24*(6200000(1+15%)/5)

[{(404(0.9)-285)*30000}-1500000]*0.76 + 0.24*(6200000(1.15))/5)

652080+342240 = $994320

NPV = -$6200000(1.15)-$600000(1.05)+$994320(PVIFA13%,5 ) + [$600000(1.05)+$875000(0.85)(1-0.24)]/1.135

-$6200000(1.15)-$600000(1.05)+$994320(3.5172) + [$600000(1.05)+$875000(0.85)(1-0.24)]/1.135

-7130000-630000+3497222.304+((630000 +565250)/1.8424)

NPV = -$3614031.50

Best case:

OCF

[{(404(1.1)-285)*30000}-1500000]*0.76 + 0.24*(6200000(0.85))/5)

2494320+252960 = $2747280

NPV

-$6200000(0.85)-$600000(0.95)+$2747280(3.5172) + [$600000(0.95)+$875000(1.15)(1-0.24)]/1.135

-5270000-570000+9662733.216+ [570000+ 764750]/1.8424

= $4547195.86